# Polytope Volume Monitoring Problem: Formulation and Solution via Parametric Linear Program Based Control Barrier Function

### v1

@@ III-B Preliminary Analysis
The center $c\in\mathbb{R}^{l}$ and radius $r\geq 0$ of the largest sphere contained inside the polytope $\Phi(x)$ defined in (7b), called Chebyshev ball [21, Chapter 8], can be computed by the following Linear Program at every $x$:
$r^{*}(x):=$ $\max_{c,r}~{}r=-\min_{c,r}~{}-r,$ (11a)
s.t. $a_{j}^{\top}(x)c+\|a_{j}(x)\|r\leq b_{j}(x),~{}\forall j\in\mathcal{J},$ (11b)
then from the geometric meaning in Fig. 1(a), it holds that
$\mathcal{V}(x(t))\geq\pi(r^{*}(x(t)))^{2}\geq\epsilon,$ (12)
from which one knows the control goal in (8) would hold if
$r^{*}(x(t))\geq\epsilon_{0}:=\sqrt{\frac{\epsilon}{\pi}},~{}\forall t\geq 0.$ (13)
@@ IV Main results
To solve the aforementioned open problem, the nonsmooth parametric optimization theory, as shown in Appendix -A, is introduced to develop a general approach for synthesizing a NCBF-based safety filter to solve PVM problems for the first time. Without calculating the generalized gradient $\partial r^{*}(x)$, the NCBF condition is established based on its directional derivative for convenience.
@@ IV-A Establishing a Parameteric LP-defined NBF Candidate
Inspired by (13), one can define an implicit function $h_{r}(x)$ via the following parametric linear program:
$h_{r}(x):=$ $r^{*}(x)-\epsilon_{0}=-\epsilon_{0}-\min_{z\in F_{P}(x)}~{}q^{\top}z,$ (14)
where $q:=\operatorname{col}{(0_{l\times 1},-1)}$ and $z:=\operatorname{col}{(c,r)}\in\mathbb{R}^{l+1}$. The feasible set $F_{P}(x)$ is defined as
$F_{P}(x):=\left\{z\mid\tilde{\mathbf{A}}(x)z\leq\tilde{\mathbf{b}}(x)\right\},$ (15)
where $\tilde{\mathbf{A}}(x)\in\mathbb{R}^{N\times(l+1)}$ and $\tilde{\mathbf{b}}(x)\in\mathbb{R}^{N}$ are
$\tilde{\mathbf{A}}(x):=\left[\begin{array}[]{cc}\tilde{a}_{1}(x)^{\top}\\ \tilde{a}_{2}(x)^{\top}\\ \cdots\cdots\\ \tilde{a}_{N}(x)^{\top}\end{array}\right],\quad\tilde{\mathbf{b}}(x):=\left[\begin{array}[]{c}b_{1}(x)\\ b_{2}(x)\\ \cdots\cdots\\ b_{N}(x)\end{array}\right],$ (24)
with $\tilde{a}_{j}(x):=\operatorname{col}(a_{j}(x),\|a_{j}(x)\|)$ for every $j=1,2,...,N$.
Next, preliminary properties of the PLP in (14) are shown first, where Mangasarian Fromovitz Constraint Qualification (MFCQ), as in Definition 3, is a key condition in studying the continuity and directional differentiability of $r^{*}(x)$.
@@
For the value function $h_{r}(x)$ of the PLP in (14), the following conclusions hold:
@@
    The feasible set $F_{P}(x)$ is nonempty for every $x$ and uniformly bounded for all $x$.
@@
    For every $x$, $h_{r}(x)$ is finite, and the solution map
    $S(x):=$ $\operatorname*{argmin}_{z\in F_{P}(x)}~{}q^{\top}z,$ (25)
@@
    For every $x$, the MFCQ holds at any $z\in F_{P}(x)$.
@@
To avoid breaking the flow of the exposition, please see the note in Appendix -B. ∎
Based on Proposition 1, one can prove that $h_{r}(x)$ can be chosen as a NBF candidate in the sense of (2).
@@
The Parametric LP defined implicit function $h_{r}(x)$ in (14) is locally Lipschitz continuous, and it is a NBF candidate on closed set $\mathcal{S}:=\{x:h_{r}(x)\geq 0\}$.
@@
Recalling Proposition 1 i) and iii) that ${F}_{P}(x)$ is uniformly bounded for all $x$ and the MFCQ condition holds, according to Lemma 2, $h_{r}(x)$ is locally Lipschitz continuous at every $x$. From (14), the conditions in (2) hold for $h_{r}(x)$, hence $h_{r}(x)$ is a NBF candidate on $\mathcal{S}$. ∎
Before going to develop the theoretical results, the geometric structure of the solution map $S(x)$ is discussed, which is helpful to understand the expression of $h_{r}^{\prime}(x;\dot{x})$ in (28). First, from the definition (25), $S(x)$ can be rewritten as
$S(x)$ $=T(x)\times\left\{r\mid r=r^{*}(x)\right\},$
where
$T(x):=\bigcap_{j\in\mathcal{J}}\left\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x)-\|a_{j}(x)\|r^{*}(x)\right\}.$
Since $r^{*}(x)\geq 0$, one has
$T(x)\subseteq\left\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x),\forall j\in\mathcal{J}\right\}=\Phi(x)\subset\mathcal{D}_{y}.$
Because $T(x)$ is expressed as the intersection of halfspcaes, it is still a (convex and compact) polytope in $\mathcal{D}_{y}\subset\mathbb{R}^{l}$, whose dimension, lower than $\Phi(x)$, may be $10,1,2,...,l-$ 1. In the above, dimension 00 means that $T(x)$ reduces to a singleton.
@@ IV-B Building the NCBF Condition
This section aims to establish the NCBF condition based on existing results about the directional differentiability of parametric convex programs in Lemma 3.
In detail, recalling the parameteric linear program in (14):
$v_{r}(x):=\min_{z\in F_{P}(x)}~{}q^{\top}z=-h_{r}(x)-\epsilon_{0},$
then the directional derivatives of $v_{r}(x)$ and $h_{r}(x)$ are studied.
First, according to the feasible set $F_{P}(\cdot)$ in (15), the active index set [^1] Inspired by [19], $\mathcal{J}_{0}({z},x)$ can be replaced by the socalled almost-active index set $\mathcal{J}_{\varepsilon}({z},x):=\{j\in\mathcal{J}\mid\|\tilde{a}_{j}(x)^{\top}{z}-b_{j}(x)\|\leq\varepsilon\}$ for real implementation, where $\varepsilon$ is a sufficiently small scalar, such as $\varepsilon=10^{-4}$. is defined for every ${x}$ and feasible ${z}\in F_{P}({x})$ as:
$\mathcal{J}_{0}({z},x):=$$\left\{j\in\mathcal{J}\mid\tilde{a}_{j}(x)^{\top}{z}=b_{j}(x)\right\}$ (26)
$=$ $\left\{j\in\mathcal{J}\mid a_{j}^{\top}(x)c+\|a_{j}(x)\|r=b_{j}(x)\right\},$
where $\tilde{a}_{j}(x):=\operatorname{col}(a_{j}(x),\|a_{j}(x)\|)$.
For every $\bar{z},\bar{x}$ satisfying $\bar{z}\in S(\bar{x})$, recalling $S(\bar{x})\subset\partial F_{P}(\bar{x})$, it holds that
@@
Based on (24), one can further define the active matrix of $\tilde{\mathbf{A}}(x)$ and the active vector of $\tilde{\mathbf{b}}(x)$ for every $x$ as:
$\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x):=\left[\begin{array}[]{c}\tilde{a}_{j_{1}}(x)^{\top}\\ \tilde{a}_{j_{2}}(x)^{\top}\\ ......\\ \tilde{a}_{j_{m}}(x)^{\top}\end{array}\right],\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x):=\left[\begin{array}[]{c}b_{j_{1}}(x)\\ b_{j_{2}}(x)\\ ......\\ b_{j_{m}}(x)\end{array}\right],$ (27)
where $j_{1},j_{2},...,j_{m}$ represent all indices in $\mathcal{J}_{0}(\bar{z},\bar{x})$. In the above, variables “$\bar{z},\bar{x}$”, appearing in the active set $\mathcal{J}_{0}(\cdot,\cdot)$, are introduced to make a difference from the variable “$x$” in $\tilde{a}_{j}(x)$ and $b_{j}(x)$. Then the directional derivative of $h_{r}(x)$ along the system can be obtained through Lemma 3.
@@
The directional derivative of $h_{r}(x)$ in (14) at $x$ along $\dot{x}$ can be expressed as
@@
$\partial_{x}L(\mu,z,x)$ $=\sum_{j\in\mathcal{J}_{0}({z},{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial{{b}}_{j}}{\partial x}(x)\right),$ (29)
and the dual feasible set
@@
To avoid breaking the flow of the exposition, please see the note in Appendix -B. ∎
@@
$\max_{\begin{subarray}{c}z\in S(x)\\ u\in\Psi(x)\end{subarray}}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\left(f+gu\right)\geq-\alpha(h_{r}(x)).$ (31)
@@
The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (4) holds for $h_{r}(\cdot)$, if there exist $\alpha,\epsilon>0$, for all $x\in\mathcal{D}_{\epsilon}$, $\exists u\in\Psi(x)$, such that
@@
where $z^{*}(x)=\operatorname{col}{(c^{*}(x),r^{*}(x))}$ is chosen from $S(x)$ and the expression of $\partial_{x}L(\cdot)$ is given in (29).
@@
$\geq~{}\min_{\mu\in M(z^{*}(x),x)}-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$
$\geq-\alpha h_{r},$

### v3

@@ III-B Preliminary Analysis
The radius $r\geq 0$ of the largest sphere contained inside the polytope $\Phi(x)$ defined in (6b), called Chebyshev ball [21, Chapter 8], can be computed by the following linear program (LP) at every $x$:
$r^{*}(x)\mathrel{\mathop{\ordinarycolon}}$= $\max_{c\in\mathbb{R}^{l},r\geq 0}~r=-\min_{c\in\mathbb{R}^{l},r\geq 0}~-r,$ (10a)
s.t. $a_{j}^{\top}(x)c+\|a_{j}(x)\|r\leq b_{j}(x),~\forall j\in\mathcal{J}.$ (10b)
Given any user-defined threshold value $\epsilon_{0}>0$, as illustrated in Fig. 1 (a), one can know that the goal in (7) holds if
$r^{*}(x(t))\geq\epsilon_{0},~\forall t\geq 0.$ (11)
@@ IV Main results
To solve the aforementioned open problem, the nonsmooth parametric optimization theory [18], as shown in Appendix A-A, is introduced to develop a general approach for synthesizing a NCBF-based safety filter to solve PVM problems for the first time. Without calculating the generalized gradient $\partial r^{*}(x)$, the NCBF condition is established based on its directional derivative for convenience.
@@ IV-A Establishing a Parameteric LP-defined NBF Candidate
Inspired by (10) and (11), one can define an implicit function $h_{r}(x)$ via the following parametric linear program:
$h_{r}(x)\mathrel{\mathop{\ordinarycolon}}$= $r^{*}(x)-\epsilon_{0}=-\epsilon_{0}-\!\min_{z\in F_{P}(x)}q^{\top}z,\forall x\in\mathcal{A},$ (12)
where $q\mathrel{\mathop{\ordinarycolon}}=\operatorname{col}{(0_{l\times 1},-1)}$ and $z\mathrel{\mathop{\ordinarycolon}}=\operatorname{col}{(c,r)}\in\mathbb{R}^{l+1}$. The feasible set $F_{P}(x)$ is defined as
$F_{P}(x)\mathrel{\mathop{\ordinarycolon}}=\left\{z\mid\tilde{\mathbf{A}}(x)z\leq\tilde{\mathbf{b}}(x)\right\},$ (13)
where $\tilde{\mathbf{A}}(x)\in\mathbb{R}^{(N+1)\times(l+1)}$ and $\tilde{\mathbf{b}}(x)\in\mathbb{R}^{N+1}$ are
$\tilde{\mathbf{A}}(x)\mathrel{\mathop{\ordinarycolon}}=\left[\!\!\!\begin{array}[]{cc}0_{1\times l},&-1\\ a_{1}^{\top}(x),&\|a_{1}(x)\|\\ a_{2}^{\top}(x),&\|a_{2}(x)\|\\ \cdots&\cdots\\ a_{N}^{\top}(x),&\|a_{N}(x)\|\end{array}\!\!\!\right],~\tilde{\mathbf{b}}(x)\mathrel{\mathop{\ordinarycolon}}=\left[\!\!\!\begin{array}[]{c}0\\ b_{1}(x)\\ b_{2}(x)\\ \cdots\\ b_{N}(x)\end{array}\!\!\!\right].$ (24)
For brevity, let $\tilde{a}_{j}^{\top}(x)$ and $\tilde{b}_{j}(x)$ denote the $j$-th row of $\tilde{\mathbf{A}}(x)$ and $\tilde{\mathbf{b}}(x)$ respectively, i.e.,
$\tilde{a}_{j}(x)\mathrel{\mathop{\ordinarycolon}}=\tilde{\mathbf{A}}^{\top}(x)e_{j},~\tilde{b}_{j}(x)\mathrel{\mathop{\ordinarycolon}}=\tilde{\mathbf{b}}^{\top}(x)e_{j},\forall j\in[N+1],$
where $e_{j}$ is the $j$-th column of the $(N+1)$-dimensional identity matrix and $N+1]\mathrel{\mathop{\ordinarycolon}}=\{1,2,\cdots,N+1\}$.
Next, the basic properties of the parametric linear program in (12) are shown first, where Mangasarian Fromovitz Constraint Qualification (MFCQ), as in Definition A.1, is a key condition in studying the continuity and directional differentiability of $r^{*}(x)$.
@@
Consider the set $\mathcal{A}$ defined in (7) and the function $h_{r}(x)$ in (12). The following conclusions hold:
@@
    The interior of $F_{P}(x)$ is nonempty for all $x\in\mathcal{A}$.
@@
    Furthermore, the set $F_{P}(x)$ and the function $h_{r}(x)$ are uniformly bounded for all $x\in\mathcal{A}$. The solution map
    $S(x)\mathrel{\mathop{\ordinarycolon}}$= $\operatorname*{argmin}_{z\in F_{P}(x)}~q^{\top}z,$ (25)
@@
    For all $x\in\mathcal{A}$, the MFCQ holds at any $z\in F_{P}(x)$.
@@
i) First, it is shown that $F_{P}(x)\neq\emptyset,\forall x\in\mathcal{A}$. Recalling
$\operatorname{Int}(\Phi(x))=\left\{c\mid a_{j}^{\top}(x)c<b_{j}(x),~\forall j\in\mathcal{J}\right\},$
for $x\in\mathcal{A}$, since $\operatorname{Int}(\Phi(x))$ is nonempty and uniformly bounded (according to Assumption 1), one can find a sufficiently small $\varepsilon_{0}(x)>0$ and a point $c_{0}(x)\in\mathcal{D}_{y}$, such that
$a_{j}^{\top}(x)c_{0}(x)+\varepsilon_{0}(x)<b_{j}(x),~\forall j\in\mathcal{J}.$
For every $j$, since $\|a_{j}(x)\|\neq 0$ as in Assumption 1, there exists $c_{2}(x)\neq 0$, s.t.
$\varepsilon_{0}=a_{j}^{\top}c_{2}=\|a_{j}\|\underbrace{\|c_{2}\|\cos\langle a_{j},c_{2}\rangle}_{=\mathrel{\mathop{\ordinarycolon}}r_{j}(x)}>0.$
Let $r_{0}\mathrel{\mathop{\ordinarycolon}}=\min_{j\in\mathcal{J}}~r_{j}$, one has $r_{0}>0$, $\|a_{j}\|r_{0}\leq\varepsilon_{0}$, and
$a_{j}^{\top}(x)c_{0}(x)+\|a_{j}(x)\|r_{0}(x)<b_{j}(x),~\forall j\in\mathcal{J},$
which means that $\operatorname{Int}(F_{P}(x))\neq\emptyset$.
ii) The following shows the uniform boundedness of $F_{P}(x)$. Recalling the expression of $F_{P}(x)$ in (13):
$F_{P}(x)\mathrel{\mathop{\ordinarycolon}}=\left\{z\mid r\geq 0,{a}_{j}^{\top}c+\|{a}_{j}\|r\leq{b}_{j},\forall j\in\mathcal{J}\right\},$
from which one knows that
$0\leq r\leq$ $\max_{c\in\Phi(x)}\min_{j\in\mathcal{J}}\frac{{b}_{j}(x)-{a}_{j}^{\top}(x)c}{\|{a}_{j}(x)\|}$
$=$ $\max_{c\in\Phi(x)}d(c,\partial\Phi(x))\leq\max_{c\in\mathcal{D}_{y}}d(c,\partial\mathcal{D}_{y})=\mathrel{\mathop{\ordinarycolon}}r_{d},$
where $d(c,\mathcal{C})\mathrel{\mathop{\ordinarycolon}}=\min_{c^{\prime}\in\mathcal{C}}\|c-c^{\prime}\|$ denotes the distance function from a point $x$ and a set $\mathcal{C}$, the compact set $\mathcal{D}_{y}$ is derived from Assumption 1 and $r_{d}$ is a positive constant. Furthermore, one has
$F_{P}(x)\subset\Phi(x)\times[0,r_{d}]\subset\mathcal{D}_{y}\times[0,r_{d}],$
which means that the set $F_{P}(x)$ is uniformly bounded, and functions $r^{*}(x)$ and $h_{r}(x)$ are also uniformly bounded. Due to the fact that (12) is a LP problem, the optimization value is obtained at a point $z$ only if $z$ is on the boundary $\partial F_{P}(x)$.
iii) The MFCQ condition holds for the optimization problem in (12) since it is a LP piecewise at $x$. Recalling the fact in [23] that for $\mathcal{C}^{1}$ convex programs with affine inequalities, the MFCQ condition is equivalent to the Slater condition. Since we have proved in Theorem 1 i) that $\operatorname{Int}(F_{P}(x))\neq\emptyset,\forall x\in\mathcal{A}$, there exists a point $z\in\operatorname{Int}(F_{P}(x))$ such that the inequalities in (13) holds strictly, hence the Slater condition (as well as the MFCQ condition) naturally holds at any $z\in F_{P}(x)$. ∎
Based on Proposition 1, one can prove that $h_{r}(x)$ is a NBF candidate by verifying the conditions in Definition 1.
@@
The Parametric LP defined implicit function $h_{r}(x)$ in (12) is locally Lipschitz continuous on $\mathcal{A}$, and it is a NBF candidate on the closed set $\mathcal{S}\mathrel{\mathop{\ordinarycolon}}=\{x\mid h_{r}(x)\geq 0\}\subset\mathcal{A}$.
@@
Recalling Proposition 1 ii) and iii) that ${F}_{P}(x)$ is uniformly bounded for $x\in\mathcal{A}$ and the MFCQ condition holds, the compact set $C_{0}$ required in Lemma A.1 exists, hence the conditions in Lemma A.1 are satisfied, thus $h_{r}(x)$ is locally Lipschitz continuous. Since $\mathcal{S}$ is defined by $h_{r}(x)$, $h_{r}(x)$ is a NBF candidate on the set $\mathcal{S}$ according to Definition 1. ∎
@@ IV-B Building the NCBF Condition
This section aims to establish the NCBF condition based on existing results about the directional differentiability of parametric convex programs in Lemma A.2.
First, according to the feasible set $F_{P}(\cdot)$ in (13), the active index set [^1] Inspired by [19], $\mathcal{J}_{0}({z},x)$ can be replaced by the socalled almost-active index set $\mathcal{J}_{\varepsilon}({z},x)\mathrel{\mathop{\ordinarycolon}}=\{j\in\mathcal{J}\mid\|\tilde{a}_{j}^{\top}(x){z}-\tilde{b}_{j}(x)\|\leq\varepsilon\}$ for real implementation, where $\varepsilon$ is a sufficiently small scalar, such as $\varepsilon=10^{-4}$. is defined for every x{x} and feasible ${z}\in F_{P}({x})$ as:
$\mathcal{J}_{0}({z},x)\mathrel{\mathop{\ordinarycolon}}$= $\left\{j\in[N+1]\mid\tilde{a}_{j}^{\top}(x){z}=\tilde{b}_{j}(x)\right\}.$ (26)
For every $\bar{z},\bar{x}$ satisfying $\bar{z}\in S(\bar{x})$, recalling $S(\bar{x})\subset\partial F_{P}(\bar{x})$ in Proposition 1 iii), it holds that
@@
Based on (24) and (26), one can further define the active matrix of $\tilde{\mathbf{A}}(x)$ and the active vector of $\tilde{\mathbf{b}}(x)$ for every $x$:
$\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)\mathrel{\mathop{\ordinarycolon}}=\!\!\left[\!\!\!\begin{array}[]{c}\tilde{a}_{j_{1}}^{\top}(x)\\ \tilde{a}_{j_{2}}^{\top}(x)\\ ......\\ \tilde{a}_{j_{m}}^{\top}(x)\end{array}\!\!\!\right],\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)\mathrel{\mathop{\ordinarycolon}}=\!\!\left[\!\!\!\begin{array}[]{c}\tilde{b}_{j_{1}}(x)\\ \tilde{b}_{j_{2}}(x)\\ ......\\ \tilde{b}_{j_{m}}(x)\end{array}\!\!\!\right],$ (27)
where $j_{1},j_{2},...,j_{m}$ represent all indices in $\mathcal{J}_{0}(\bar{z},\bar{x})$. In the above, variables $\bar{z},\bar{x}$, appearing in the active set $\mathcal{J}_{0}(\cdot,\cdot)$, are introduced to make a difference from the variable $x$ in $\tilde{a}_{j}(x)$ and $\tilde{b}_{j}(x)$. Then the directional derivative of $h_{r}(x)$ along the system can be obtained through Lemma A.2.
@@
The directional derivative of $h_{r}(x)$ in (12) at $x$ along $\dot{x}$ is
@@
$\partial_{x}L(\mu,z,x)$ $=\!\sum_{j\in\mathcal{J}_{0}({z},{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial\tilde{b}_{j}}{\partial x}(x)\right),$ (29)
and the set of multipliers
@@
This proposition is the corollary of Lemma A.2 with the boundness of set $M(z,x)$ to be proven. To prove this, first define
$v_{r}(x)\mathrel{\mathop{\ordinarycolon}}=\min_{z\in F_{P}(x)}~q^{\top}z,$
then $h_{r}(x)$ can be expressed as $h_{r}(x)=-v_{r}(x)-\epsilon_{0}$.
According to (A.3), the local Lagrangian function near $\bar{x},\bar{z}$ is defined as:
$=\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial\tilde{{b}}_{j}}{\partial x}(x)\right),$
and $M(z,x)\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\partial_{z}L(\mu,z,x)=0^{\top},~\mu\geq 0\right\}$. In the above, the partial derivative of $L(\mu,z,x)$ w.r.t. variable $z,x$ at the point $\bar{z},\bar{x}$ are calculated. By further choosing $\bar{z}=z$ and $\bar{x}=x$, the expressions in (29) and (30) can be obtained.
Since MFCQ condition holds and $F_{P}(x)$ is compact as proved in Proposition 1, then the conditions of Lemma A.2 hold, from which one has
$h^{\prime}_{r}(x;\dot{x})$ $=-v^{\prime}_{r}(x;\dot{x})=-\min_{z\in S(x)}\max_{\mu\in M(z,x)}\partial_{x}L(\mu,z,x)\dot{x}$
$=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\dot{x}.$
Hence, the expression in (28) is obtained.
The set $M(z,x)$ in (30) can be further rewritten as:
$M(z,x)$ $=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\tilde{a}_{j}(x)\mu_{j}=-q,~\mu\geq 0\right\}$
$=M_{1}(z,x)\cap M_{2}(z,x),$
where
$M_{1}(z,x)$ $\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}a_{j}(x)\mu_{j}=0\right\},$
$M_{2}(z,x)$ $\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\|a_{j}(x)\|\mu_{j}=1,~\mu\geq 0\right\}.$
From the structure of $M_{2}(z,x)$, since $\|a_{j}(x)\|>0$ from Assumption 1, $M_{2}(z,x)$ is a bounded cone [21], hence its subset $M(z,x)$ is bounded naturally. ∎
@@
$\!\max_{\begin{subarray}{c}z\in S(x)\\ u\in\Psi(x)\end{subarray}}\min_{\mu\in M(z,x)}\!\!-\partial_{x}L(\mu,z,x)\left(f+gu\right)\geq-\alpha(h_{r}).$ (31)
@@
The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (3) holds for $h_{r}(\cdot)$, if there exist $\alpha,\bar{\epsilon}>0$, for all $x\in\mathcal{D}_{\bar{\epsilon}}\mathrel{\mathop{\ordinarycolon}}=\{x\mid h_{r}(x)\geq-\bar{\epsilon}\}$, $\exists u\in\Psi(x)$, such that
@@
where $z^{*}(x)=\operatorname{col}{(c^{*}(x),r^{*}(x))}$ is any point chosen from the set $S(x)$ and the expression of $\partial_{x}L(\cdot)$ is given by (29).
@@
$\geq~\min_{\mu\in M(z^{*}(x),x)}-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)\geq-\alpha h_{r},$